WARNING: This web is not used anymore. Please use PDBService.MinuteS28February08 instead!
 

Minutes 28 February 08

  • Phone:
    • Gridka: Doris
    • RAL: Carmine
    • NDGF: Olli
    • BNL: Carlos
    • CNAF: Barbara
    • LHCb: Marco Clemencic

  • CERN:
    • 3D: Dirk, Eva
    • ATLAS: Gancho, Florbela

  • General announcements:
    • Agenda for the next workshop. Dates not completely confirmed, but probably in the week from 7th to 11th of April. Most likely in Lyon. Confirmation next week.
    • Thanks to Gordon for his feedback on the FTS database. Miguel Anjo will send a summary for Tier0.

  • Status Tier0 and Tier1 sites:
    • Tier0: CPU patch, Streams bugs patches and logminer tables maintenance.
      • CCRC running smoothly. No high impact at Tier0.
      • Dirk still working on the licenses agreement with Oracle.
    • RAL: CPU patch + recompilation to be done in March.
      • TAG database almost ready, some last tests to be performed.
    • NDGF: new hw progressing slowly
    • BNL: CPU patch applied, recompilation pending, probably next week, announcement will be sent
      • Planning to get new storage.
      • Problem getting mails from the 3d interventions panel. Please let us know if you insert new intervention and the email is not sent.
      • Thanks to Eva for adding the information about sessions and connections coming from last ATLAS tests.
    • GridKA: nothing to report
    • CNAF: CPU patch on the LHCb cluster, no views replication, postponed to March, CPU patch also for other databases.
      • ATLAS conditions cluster, recommendations implemented.
      • Big intervention, last week of March, complete downtime during 7 days, moving computer centre.
      • VOMS functionality tests done at CNAF and between CERN and CNAF. Now preparing to run with more than one VO.

  • Experiment status:
    • LHCb: nothing to report
    • ATLAS:
      • DQ2 – Streams replication required, in particular to SARA. Streams setup between CERN and SARA, but apply process misbehaving last Saturday and it was dropped. * This setup will move to test machines because the application is still being developed. Discussion about different ways of disaster recovery for this application ongoing. * Contact persons: Pedro Salgado and Miguel Branco. * Dirk proposed a meeting.
      • TAGs request on streaming some small part of data, but this is a plan on long term.
      • AMI replication from Lyon to CERN – topic for the next workshop?
      • MUON sites replication to CERN: 3 sites to be joined. First run some tests on the validation environment.
        • Responsible names for the MUON sites databases needed.
        • Gancho is helping them to setup the databases.
      • PVSS replica move to production in April.

  • Workshop agenda:
    • server limits, connection limits, sniping – Carlos and Andrew
    • 32 bits to 64 bits migration in place - Carlos
    • new storage systems installed – Luis and Luca
    • typical interventions and what is doing by whom – Eva and Maria, Barbara
    • experiment connectivity request: who is accessing the databases, firewall configurations – ?
    • transportable tablespaces and reinstantiation after long outages – Eva
    • experiment plans for next CCRC
    • Lyon and AMI
    • FTS and LFC installation – Gordon
    • Dirk will send around the agenda together with the confirmed date for the workshop.

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2008-04-21 - EvaDafonte
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback